OpenStudy (anonymous):

hello i really need help with some questions.... A rifle is aimed horizontally toward the center of a target 100m away. if the bullet strikes 10cm below the center, what was the velocity of the bullet? (ignore air friction.)

OpenStudy (amistre64):

you should have equations of motion to play with if you do not have calculus to do your work for you.

OpenStudy (amistre64):

h = gt^2/2 solving for t, 2h/g = t^2 sqrt(2h/g) = t is how long it takes to 'drop' a certain height. velocity is just distance/time ....